# An extension of May’s Theorem to three alternatives: axiomatizing Minimax voting

### v2

@@ Abstract
May’s Theorem [K. O. May, Econometrica 20 (1952) 680-684] characterizes majority voting on two alternatives as the unique preferential voting method satisfying several simple axioms. Here we show that by adding some desirable axioms to May’s axioms, we can uniquely determine how to vote on three alternatives (setting aside tiebreaking). In particular, we add two axioms stating that the voting method should mitigate spoiler effects and avoid the so-called strong no show paradox. We prove a theorem stating that any preferential voting method satisfying our enlarged set of axioms, which includes some weak homogeneity and preservation axioms, must choose from among the Minimax winners in all three-alternative elections. When applied to more than three alternatives, our axioms also distinguish Minimax from other known voting methods that coincide with or refine Minimax for three alternatives.
@@ 1 Introduction
The voting method determined by our axioms (modulo tiebreaking) is the Minimax method (Simpson 1969, Kramer 1977), which chooses the alternative that minimizes the maximal loss, if any, in head-to-head majority comparisons with other alternatives. In two-alternative elections, Minimax reduces to majority voting. In three-alternative elections, it apparently coincides with Condorcet’s (1785) own approach to resolving cycles with three alternatives (Young 1988, p. 1233). In fact, a number of sophisticated Condorcet-consistent voting methods coincide with or refine Minimax in three-alternative elections: Kemeny (Kemeny 1959), Ranked Pairs (Tideman 1987), Beat Path (Schulze 2011), Split Cycle (Holliday and Pacuit 2023a), and Stable Voting Holliday and Pacuit 2023b). Thus, our result provides support for this common solution to the problem of voting on three alternatives. Of course, different sets of axioms lead to characterizations of other voting methods; [^1] See, e.g., Young 1974 and Nitzan and Rubinstein 1981 on the Borda method, Richelson 1978, Ching 1996, and Sekiguchi 2012 on the Plurality method, Merlin 2003 on scoring methods, Henriet 1985 on the Copeland method, Freeman et al. 2014 on the Instant Runoff method, and Holliday and Pacuit 2021a and Ding et al. Forthcoming on the Split Cycle method. some of these characterizations may also be considered extensions of May’s Theorem (e.g., Maskin 2024); and we leave for future discussion the normative question of which set of axioms is preferable.
While our main goal is to identify axioms characterizing the common solution to voting on three alternatives given by the Condorcet methods mentioned above, we will state axioms that can also be applied to any number of alternatives. Significantly, for more than three alternatives, our axioms distinguish Minimax from all other known voting rules. Ultimately all of the Condorcet methods listed above should be axiomatically characterized with no restriction on the number of alternatives, as has been recently done for Split Cycle (Ding et al. Forthcoming). Our axioms may be a step in this direction for Minimax.
@@ 2 Preliminaries
Fix sets $\mathcal{V}$ and $\mathcal{X}$ of voters and alternatives, respectively. Assume $\mathcal{V}$ is countably infinite and $|\mathcal{X}|\geq 3$. Given nonempty finite subsets $V\subseteq\mathcal{V}$ and $X\subseteq\mathcal{X}$, a $(V,X)$-profile is a function $\mathbf{P}$ assigning to each $i\in V$ a strict weak order [^2] A strict weak order on a set $X$ is a binary relation $P$ that is asymmetric (if $(a,b)\in P$, then $(b,a)\not\in P$) and negatively transitive (if $(a,b)\not\in P$ and $(b,c)\not\in P$, then $(a,c)\not\in P$). It follows that $P$ is transitive. A strict weak order is simply the strict part of a complete and transitive relation, i.e., a ranking allowing ties. However, we will not actually need the binary relations $\mathbf{P}(i)$ to satisfy any of these properties for our results to hold, as explained in Remark 3.9. $\mathbf{P}(i)$ on $X$. If $(a,b)\in\mathbf{P}(i)$, we say that $i$ ranks $a$ above $b$; if $(a,b)\not\in\mathbf{P}(i)$ and $(b,a)\not\in\mathbf{P}(i)$, we say that $i$ ranks $a$ and $b$ as tied. We write $V(\mathbf{P})$ for the relevant set $V$ of voters and $X(\mathbf{P})$ for the relevant set $X$ of alternatives. A profile is a $(V,X)$-profile for some nonempty finite $V\subseteq\mathcal{V}$ and $X\subseteq\mathcal{X}$. Given a profile $\mathbf{P}$ and set $Y\subsetneq X(\mathbf{P})$ of alternatives, let $\mathbf{P}_{\mid Y}$ be the profile obtained by restricting each $\mathbf{P}(i)$ to $Y$. Given $b\in X(\mathbf{P})$, let $\mathbf{P}_{-b}=\mathbf{P}_{\mid X(\mathbf{P})\setminus\{b\}}$.
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ in some set $\mathrm{dom}(F)$ of profiles a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$. For example, according to the method $Maj$ of majority voting on the domain of two-alternative profiles, for any such profile $\mathbf{P}$ and alternatives $a,b\in X(\mathbf{P})$, we have $a\in Maj(\mathbf{P})$ if and only if in $\mathbf{P}$ at least as many voters rank $a$ above $b$ as rank $b$ above $a$. To state May’s Theorem characterizing this voting method, we need the three axioms on voting methods in Definition 2.1. Informally, anonymity says that the names of voters do not matter (so all voters are treated equally); neutrality says that the names of alternatives do not matter (so all alternatives are treated equally); and positive responsiveness says that improving a winning alternative’s place in a voter’s ranking not only keeps that alternative a winner [^3] As required by so-called non-negative responsiveness (see, e.g., Tideman 1987). but also makes that alternative the unique winner (so ties are easily broken).
@@
    $F$ satisfies anonymity if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ with $X(\mathbf{P})=X(\mathbf{P}^{\prime})$, if there is a permutation $\pi$ of $\mathcal{V}$ such that $\pi[V(\mathbf{P})]=V^{\prime}(\mathbf{P})$ and $\mathbf{P}(i)=\mathbf{P}^{\prime}(\pi(i))$ for all $i\in V(\mathbf{P})$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
@@
    $F$ satisfies weak positive responsiveness if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $a\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter in $V(\mathbf{P})$ who ranked $a$ uniquely last in $\mathbf{P}$ switching to ranking $a$ uniquely first in $\mathbf{P}^{\prime}$, otherwise keeping their ranking the same, then $F(\mathbf{P}^{\prime})=\{a\}$.
@@
We add the adjective ‘weak’ in ‘weak positive responsiveness’ because of our assumption that the relevant voter ranked alternative $a$ uniquely last and switches to ranking $a$ uniquely first. This makes our statement weaker than May’s (1952) original statement of positive responsiveness in two-alternative elections, which implies that if $a\in F(\mathbf{P})$ and the voter ranked $a$ and $b$ as tied, then their moving $a$ above $b$ ensures $F(\mathbf{P}^{\prime})=\{a\}$, and that if the voter ranked $b$ above $a$, then their making $a$ and $b$ tied ensures $F(\mathbf{P}^{\prime})=\{a\}$. Our statement is also weaker than later generalizations of positive responsiveness to elections with more than two alternatives (see, e.g., Cato 2011, 2018), even if voters’ rankings do not contain ties, since those generalizations imply that the voter’s moving $a$ up at all in their ranking ensures $F(\mathbf{P}^{\prime})=\{a\}$, whereas we require that the voter moves $a$ from uniquely last to uniquely first, which will matter when we turn to three-alternative elections in Section 3.
@@
(Sketch) Case 1: in $\mathbf{P}$, the same number of voters rank $a$ above $b$ as rank $b$ above $a$. Then anonymity and neutrality imply $F(\mathbf{P})=\{a,b\}$, in agreement with majority voting.
Case 2: in $\mathbf{P}$, $n$ voters rank $b$ above $a$, $n+k$ (for $k>0$) rank $a$ above $b$, and the remaining $\ell$ voters rank them as tied. Suppose for contradiction that $b\in F(\mathbf{P})$. Then by anonymity and neutrality, in any profile $\mathbf{Q}$ in which $n$ voters rank $a$ above $b$, $n+k$ rank $b$ above $a$, and the remaining $\ell$ voters rank them as tied, we have $a\in F(\mathbf{Q})$. Now let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by flipping $k$ voters who rank $a$ above $b$ to rank $b$ above $a$. Then $\mathbf{P}^{\prime}$ is a $\mathbf{Q}$ as above, so $a\in F(\mathbf{P}^{\prime})$. But by repeated application of weak positive responsiveness, $F(\mathbf{P}^{\prime})=\{b\}$, so we have a contradiction. Hence $b\not\in F(\mathbf{P})$, so $F(\mathbf{P})=\{a\}$ in agreement with majority voting.∎
@@ 3 Main result
In the rest of this section, we extend May’s Theorem to the domain of three-alternative profiles: we add axioms alongside May’s in order to force a voting method $F$ to refine Minimax, i.e., to choose from among the Minimax winners, on this domain. Thus, whenever Minimax selects a unique winner, $F$ will select the same unique winner; in the unlikely event of a Minimax tie, $F$ may perform further tie-breaking. Before adding axioms alongside May’s, we stress the importance of weak positive responsiveness in what follows: all the other axioms in our main theorem are satisfied by the trivial voting method that selects all alternatives in every profile. Thus, weak positive responsiveness is what will give our other axioms their “bite.”
@@
    $F$ satisfies positive involvement if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $a\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one new voter who ranks $a$ uniquely first, then $a\in F(\mathbf{P}^{\prime})$.
@@
Positive involvement (Saari 1995) is a variable-voter axiom: adding to an election a voter who ranks $a$ uniquely first should not cause $a$ to go from winning to losing. Pérez (2001) dubs violations of positive involvement instances of the “Strong No Show Paradox.” Surprisingly, many well-known voting methods violate positive involvement (see Pérez 2001 and Holliday and Pacuit 2021b), which Pérez calls “a common flaw in Condorcet voting correspondences.” Recall that a voting method satisfies Condorcet consistency (or the Condorcet winner criterion) if it always picks the Condorcet winner uniquely when one exists. Well-known Condorcet-consistent methods that violate positive involvement already for three alternatives include Black’s (1958) method (if there is no Condorcet winner, pick the Borda winner (Borda 1781)), Condorcet-Plurality (if there is no Condorcet winner, pick the alternative with the most first-place votes; see Lemma 4.5), Condorcet-Hare (if there is no Condorcet winner, pick the winner according to the Hare method (Hare 1859), also known as Instant Runoff Voting), and more (e.g., Baldwin (Baldwin 1926), the version of Nanson (Nanson 1882) that iteratively eliminates alternative with below average Borda score (Zwicker 2016, p. 37), Raynaud (Arrow and Raynaud 1986), Condorcet-Bucklin, Condorcet-Coombs, etc.). [^4] For example violations, see https://github.com/wesholliday/may-minimax. By contrast, though Minimax is Condorcet consistent, it does not have the flaw of violating positive involvement. [^5] Pérez (1995, Lemma 3) (cf. Pérez 2001, Lemma 1), adapting an argument of Moulin (1988), also shows that if a voting method $F$ satisfies positive involvement and Condorcet consistency, then in any profiles $\mathbf{P}$ of linear ballots (without ties), $F$ must choose from among the alternatives $a$ that Holliday (2024) calls defensible: for any alternative $b$, there is an alternative $c$ such that $\mathrm{Margin}_{\mathbf{P}}(c,b)\geq\mathrm{Margin}_{\mathbf{P}}(b,a)$ (see Kasper et al. 2019 for a study of the voting method that selects all defensible alternatives). For profiles of strict weak order ballots (allowing ties), there is an analogous result if Condorcet consistency is strengthened to the condition that if there a weak Condorcet winner (an alternative with non-negative margins against every alternative), then $F$ picks from among the weak Condorcet winners (Holliday 2024, Lemma 2.2). However, there can be Minimax losers who are defensible (see the “descending case” in the proof of Theorem 3.8), so weak Condorcet consistency and positive involvement are not sufficient for our purposes here. We will use weak positive responsiveness and two axioms to come (block preservation and homogeneity) to show that defensible alternatives that are Minimax losers will not be selected.
@@
Immunity to spoilers (Holliday and Pacuit 2023a) is a variable-alternative axiom: adding to an election a losing alternative $b$ should not cause the loss of an alternative $a$ that would win without $b$ in the election and would win in a head-to-head contest with $b$ [^6] In Holliday and Pacuit 2023a, the condition in immunity to spoilers that $a$ would win in a head-to-head contest with $b$ is expressed as the condition that more voters rank $a$ above $b$ than vice versa; but in the presence of May’s axioms that imply majority rule in two-alternative elections, the two formulations of immunity to spoilers are equivalent. The near immunity version in Definition 3.2 adds the assumption that $a$ is the unique winner without $b$ in the election, which makes for a weaker axiom. [^7] Near immunity may be more plausible in cases where some distinct alternatives $a$ and $c$ have a zero margin. Then if both $a$ and $c$ beat a third alternative $b$, but $c$ beats $b$ by a larger margin, then it may be reasonable to break the tie between $a$ and $c$ in favor of $c$. This is consistent with near immunity to spoilers but not with the stronger version. For further discussion, see the last paragraph of Section 1.1 and Definition 4.8.1 in Holliday and Pacuit 2023a, as well as the Appendix. Many well-known voting methods violate (near) immunity to spoilers (see Holliday and Pacuit 2023a). Although ranked ballots are not collected in Plurality Voting elections, the 2000 Florida Presidential Election is plausibly a case of a spoiler effect for Plurality relative to voters’ actual preferences (Magee 2003, Herron and Lewis 2007). Spoiler effects for Instant Runoff Voting include the 2009 Mayoral Election in Burlington, Vermont and the 2022 Alaska Primary Special Election for U.S. Representative (see https://github.com/voting-tools/election-analysis). By contrast, Minimax renders such spoiler effects impossible.
@@
It is important to note that the use of near immunity to spoilers in our main result (Theorem 3.8) could be replaced by Condorcet consistency. We use the framing in terms of spoilers for three reasons. First, near immunity to spoilers is strictly weaker than Condorcet consistency for three alternatives (while Condorcet consistency implies near immunity to spoilers on this domain, the converse implication fails, as shown by, e.g., the trivial voting method that selects all alternatives in every profile). Second, the framing in terms of immunity to spoilers highlights the desirability of using a Condorcet method for elections with three alternatives, such as the top-three elections advocated in Foley and Maskin 2024, since mitigating spoiler effects is a significant benefit. Third, looking ahead to profiles with more than three alternatives (Section 5), Condorcet consistency does not imply near immunity to spoilers for more than three alternatives, and in fact Minimax is quite distinctive in satisfying the axiom for any number of alternatives (see Table 1).
@@
    $F$ satisfies block preservation if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $X(\mathbf{P})=X(\mathbf{P}^{\prime})$ and $\mathbf{P}^{\prime}$ contains for each linear order of $X(\mathbf{P})$ exactly one voter submitting that linear order and no other voters, then $\mathbf{P}+\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ and $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{P}^{\prime})$.

### v4

@@ Abstract
May’s Theorem [K. O. May, Econometrica 20 (1952) 680-684] characterizes majority voting on two alternatives as the unique preferential voting method satisfying several simple axioms. Here we show that by adding some desirable axioms to May’s axioms, we can uniquely determine how to vote on three alternatives (setting aside tiebreaking). In particular, we add two axioms stating that the voting method should mitigate spoiler effects and the so-called strong no show paradox. We prove a theorem stating that any preferential voting method satisfying our enlarged set of axioms, which includes some weak homogeneity and preservation axioms, must choose from among the Minimax winners in all three-alternative elections. When applied to more than three alternatives, our axioms also distinguish Minimax from other known voting methods that coincide with or refine Minimax for three alternatives.
@@ 1 Introduction
The voting method determined by our axioms (modulo tiebreaking) is the Minimax method (Simpson 1969, Kramer 1977), which chooses the alternative that minimizes the maximal loss, if any, in head-to-head majority comparisons with other alternatives. In two-alternative elections, Minimax reduces to majority voting. In three-alternative elections, it apparently coincides with Condorcet’s (1785) own approach to resolving cycles with three alternatives (Young 1988, p. 1233). In fact, a number of sophisticated Condorcet-consistent voting methods coincide with or refine Minimax in three-alternative elections: Kemeny (Kemeny 1959), Ranked Pairs (Tideman 1987), Beat Path (Schulze 2011), Split Cycle (Holliday and Pacuit 2023a), and Stable Voting (Holliday and Pacuit 2023b). Thus, our result provides support for this common solution to the problem of voting on three alternatives. Of course, different sets of axioms lead to characterizations of other voting methods; [^1] See, e.g., Young 1974 and Nitzan and Rubinstein 1981 on the Borda method, Richelson 1978, Ching 1996, and Sekiguchi 2012 on the Plurality method, Merlin 2003 on scoring methods, Henriet 1985 on the Copeland method, Freeman et al. 2014 on the Instant Runoff method, and Holliday and Pacuit 2021a and Ding et al. 2024 on the Split Cycle method. some of these characterizations may also be considered extensions of May’s Theorem (e.g., Maskin 2024); and we leave for future discussion the normative question of which set of axioms is preferable.
While our main goal is to identify axioms characterizing the common solution to voting on three alternatives given by the Condorcet methods mentioned above, we will state axioms that can also be applied to any number of alternatives. Significantly, for more than three alternatives, our axioms distinguish Minimax from all other voting rules of which we are aware. Hopefully all of the Condorcet methods listed above will eventually be axiomatically characterized with no restriction on the number of alternatives, as has been done recently for Split Cycle (Ding et al. 2024). Our axioms may be a step in this direction for Minimax.
@@ 2 Preliminaries
Fix sets $\mathcal{V}$ and $\mathcal{X}$ of voters and alternatives, respectively. Assume $\mathcal{V}$ is countably infinite and $|\mathcal{X}|\geq 3$. Given nonempty finite subsets $V\subseteq\mathcal{V}$ and $X\subseteq\mathcal{X}$, a $(V,X)$-profile is a function $\mathbf{P}$ assigning to each $i\in V$ a strict weak order [^2] A strict weak order on a set $X$ is a binary relation $P$ that is asymmetric (if $(a,b)\in P$, then $(b,a)\not\in P$) and negatively transitive (if $(a,b)\not\in P$ and $(b,c)\not\in P$, then $(a,c)\not\in P$). It follows that $P$ is transitive. A strict weak order is simply the strict part of a complete and transitive relation, i.e., a ranking allowing ties. However, we will not actually need the binary relations $\mathbf{P}(i)$ to satisfy negative transitivity or even transitivity for our results to hold, as explained in Remark 3.9. $\mathbf{P}(i)$ on $X$. If $(a,b)\in\mathbf{P}(i)$, we say that $i$ ranks $a$ above $b$; if $(a,b)\not\in\mathbf{P}(i)$ and $(b,a)\not\in\mathbf{P}(i)$, we say that $i$ ranks $a$ and $b$ as tied. We write $V(\mathbf{P})$ for the relevant set $V$ of voters and $X(\mathbf{P})$ for the relevant set $X$ of alternatives. A profile is a $(V,X)$-profile for some nonempty finite $V\subseteq\mathcal{V}$ and $X\subseteq\mathcal{X}$. Given a profile $\mathbf{P}$ and set $Y\subsetneq X(\mathbf{P})$ of alternatives, let $\mathbf{P}_{\mid Y}$ be the profile obtained by restricting each $\mathbf{P}(i)$ to $Y$. Given $b\in X(\mathbf{P})$, let $\mathbf{P}_{-b}=\mathbf{P}_{\mid X(\mathbf{P})\setminus\{b\}}$.
A voting method is a function $F$ assigning to each profile $\mathbf{P}$ in some set $\mathrm{dom}(F)$ of profiles a nonempty subset $F(\mathbf{P})\subseteq X(\mathbf{P})$. Given voting methods $F$ and $G$, we say that $F$ refines $G$ on some domain $\mathcal{D}$ of profiles if for every $\mathbf{P}\in\mathcal{D}$, we have $F(\mathbf{P})\subseteq G(\mathbf{P})$. Note that every voting method refines itself.
According to the method $Maj$ of majority voting on the domain of two-alternative profiles, for any such profile $\mathbf{P}$ and alternatives $a,b\in X(\mathbf{P})$, we have $a\in Maj(\mathbf{P})$ if and only if in $\mathbf{P}$ at least as many voters rank $a$ above $b$ as rank $b$ above $a$. To state May’s Theorem characterizing this voting method, we need the three axioms on voting methods in Definition 2.1. Informally, anonymity says that the names of voters do not matter (so all voters are treated equally); neutrality says that the names of alternatives do not matter (so all alternatives are treated equally); and positive responsiveness says that improving a winning alternative’s place in a voter’s ranking not only keeps that alternative a winner [^3] As required by so-called non-negative responsiveness (see, e.g., Tideman 1987). but also makes that alternative the unique winner (so ties are easily broken).
@@
    $F$ satisfies anonymity if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ with $X(\mathbf{P})=X(\mathbf{P}^{\prime})$, if there is a permutation $\pi$ of $\mathcal{V}$ such that $\pi[V(\mathbf{P})]=V(\mathbf{P}^{\prime})$ and $\mathbf{P}(i)=\mathbf{P}^{\prime}(\pi(i))$ for all $i\in V(\mathbf{P})$, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
@@
    $F$ satisfies weak positive responsiveness if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ and $a\in X(\mathbf{P})$, if $a\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter in $V(\mathbf{P})$ who ranked $a$ uniquely last in $\mathbf{P}$ switching to ranking $a$ uniquely first in $\mathbf{P}^{\prime}$, otherwise keeping their ranking the same, then $F(\mathbf{P}^{\prime})=\{a\}$.
@@
We add the adjective ‘weak’ in ‘weak positive responsiveness’ because of our assumption that the relevant voter ranked alternative $a$ uniquely last and switches to ranking $a$ uniquely first. This makes our statement weaker than May’s (1952) original statement of positive responsiveness in two-alternative elections, which implies that if $a\in F(\mathbf{P})$ and the voter ranked $a$ and $b$ as tied, then their moving $a$ above $b$ ensures $F(\mathbf{P}^{\prime})=\{a\}$, and that if the voter ranked $b$ above $a$, then their making $a$ and $b$ tied ensures $F(\mathbf{P}^{\prime})=\{a\}$. Our statement is also weaker than the generalization of positive responsiveness to elections with more than two alternatives in Barberà 1977, Def. 7, even if voters’ rankings do not contain ties, since that generalization states that if $a\in F(\mathbf{P})$, then a voter moving $a$ up at all in their ranking ensures $F(\mathbf{P}^{\prime})=\{a\}$, whereas we require that the voter moves $a$ from uniquely last to uniquely first, which will matter when we turn to three-alternative elections in Section 3.
@@
(Sketch) The implication from 2 to 1 is obvious. For the implication from 1 to 2, we consider two cases. Case 1: in $\mathbf{P}$, the same number of voters rank $a$ above $b$ as rank $b$ above $a$. Then anonymity and neutrality imply $F(\mathbf{P})=\{a,b\}$, in agreement with majority voting.
Case 2: in $\mathbf{P}$, $n$ voters rank $b$ above $a$, $n+k$ (for $k>0$) rank $a$ above $b$, and the remaining $\ell$ voters rank them as tied. Suppose for contradiction that $b\in F(\mathbf{P})$. Then by anonymity and neutrality, in any profile $\mathbf{Q}$ in which $n$ voters rank $a$ above $b$, $n+k$ rank $b$ above $a$, and the remaining $\ell$ voters rank them as tied, we have $a\in F(\mathbf{Q})$. Now let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by flipping $k$ voters who rank $a$ above $b$ to rank $b$ above $a$. Then $\mathbf{P}^{\prime}$ is a $\mathbf{Q}$ as above, so $a\in F(\mathbf{P}^{\prime})$. But by repeated application of weak positive responsiveness starting from $\mathbf{P}$, $F(\mathbf{P}^{\prime})=\{b\}$, so we have a contradiction. Hence $b\not\in F(\mathbf{P})$, so $F(\mathbf{P})=\{a\}$ in agreement with majority voting.∎
@@ 3 Main result
In the rest of this section, we extend May’s Theorem to the domain of three-alternative profiles: we add axioms alongside May’s in order to force a voting method $F$ to refine Minimax, i.e., to choose from among the Minimax winners, on this domain. Thus, whenever Minimax selects a unique winner, $F$ will select the same unique winner; in the unlikely (with many voters) event of a Minimax tie, $F$ may perform further tie-breaking. Before adding axioms alongside May’s, we stress the importance of weak positive responsiveness in what follows: all the other axioms in our main theorem are satisfied by the trivial voting method that selects all alternatives in every profile, as well as by other methods that are non-trivial but still produce many ties, e.g., the Weighted Covering method of Dutta and Laslier 1999 and Pérez-Fernández and De Baets 2018 (see Ding et al. 2024, Proposition 3.16). Thus, weak positive responsiveness is what will give our other axioms their “bite.”
@@
    $F$ satisfies positive involvement if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ and $a\in X(\mathbf{P})$, if $a\in F(\mathbf{P})$ and $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding one new voter who ranks $a$ uniquely first, then $a\in F(\mathbf{P}^{\prime})$.
@@
Positive involvement (Saari 1995) is a variable-voter axiom: adding to an election a voter who ranks $a$ uniquely first should not cause $a$ to go from winning to losing. Pérez (2001) dubs violations of positive involvement the “(Positive) Strong No Show Paradox.” Surprisingly, many well-known voting methods violate positive involvement (see Pérez 2001 and Holliday and Pacuit 2021b), which Pérez calls “a common flaw in Condorcet voting correspondences.” Recall that a voting method satisfies Condorcet consistency (or the Condorcet winner criterion) if it always picks the Condorcet winner uniquely when one exists. Well-known Condorcet-consistent methods that violate positive involvement already for three alternatives include Black’s (1958) method (if there is no Condorcet winner, pick the Borda winner (Borda 1781)), Condorcet-Plurality (if there is no Condorcet winner, pick the alternative with the most first-place votes; see Lemma 4.5), Condorcet-Hare (if there is no Condorcet winner, pick the winner according to the Hare method (Hare 1859), also known as Instant Runoff Voting), and more (e.g., Baldwin (Baldwin 1926), the version of Nanson (Nanson 1882) that iteratively eliminates alternatives with below average Borda score (Zwicker 2016, p. 37), Raynaud (Arrow and Raynaud 1986), Condorcet-Bucklin, Condorcet-Coombs, etc.). [^4] For examples of violations, see https://github.com/wesholliday/may-minimax. By contrast, though Minimax is Condorcet consistent, it does not have the flaw of violating positive involvement. [^5] Pérez (1995, Lemma 3) (cf. Pérez 2001, Lemma 1), adapting an argument of Moulin (1988), also shows that if a voting method $F$ satisfies positive involvement and Condorcet consistency, then in any profile $\mathbf{P}$ of linear ballots (without ties), $F$ must choose from among the alternatives $a$ that Holliday (2024) calls defensible: for any alternative $b$, there is an alternative $c$ such that $\mathrm{Margin}_{\mathbf{P}}(c,b)\geq\mathrm{Margin}_{\mathbf{P}}(b,a)$ (see Kasper et al. 2019 for a study of the voting method that selects all defensible alternatives). For profiles of strict weak order ballots (allowing ties), there is an analogous result if Condorcet consistency is strengthened to the condition that if there is a weak Condorcet winner (an alternative with non-negative margins against every alternative), then $F$ picks from among the weak Condorcet winners (Holliday 2024, Lemma 2.2). However, there can be Minimax losers who are defensible (see the “descending case” in the proof of Theorem 3.8), so weak Condorcet consistency and positive involvement are not sufficient for our purposes here. We will use weak positive responsiveness and two axioms to come (block preservation and homogeneity) to show that defensible alternatives that are Minimax losers will not be selected.
@@
Immunity to spoilers (Holliday and Pacuit 2023a) is a variable-alternative axiom: adding to an election a losing alternative $b$ should not cause the loss of an alternative $a$ that would win without $b$ in the election and would win in a head-to-head contest with $b$ [^6] In Holliday and Pacuit 2023a, the condition in immunity to spoilers that $a$ would win in a head-to-head contest with $b$ is expressed as the condition that more voters rank $a$ above $b$ than vice versa; but in the presence of May’s axioms that imply majority rule in two-alternative elections, the two formulations of immunity to spoilers are equivalent. The near immunity version in Definition 3.2 adds the assumption that $a$ is the unique winner without $b$ in the election, which makes for a weaker axiom. [^7] Near immunity may be more plausible in cases where some distinct alternatives $a$ and $c$ have a zero margin. Then if both $a$ and $c$ beat a third alternative $b$, but $c$ beats $b$ by a larger margin, then it may be reasonable to break the tie between $a$ and $c$ in favor of $c$. This is consistent with near immunity to spoilers but not with the stronger version. For further discussion, see the last paragraph of Section 1.1 and Definition 4.8.1 in Holliday and Pacuit 2023a, as well as the Appendix of this paper. Many well-known voting methods violate (near) immunity to spoilers (see Holliday and Pacuit 2023a). Although ranked ballots are not collected in Plurality Voting elections, the 2000 Florida Presidential Election is plausibly a case of a spoiler effect for Plurality relative to voters’ actual preferences (Magee 2003, Herron and Lewis 2007). Spoiler effects for Instant Runoff Voting include the 2009 Mayoral Election in Burlington, Vermont and the 2022 Alaska Primary Special Election for U.S. Representative (see https://github.com/voting-tools/election-analysis). By contrast, Minimax renders such spoiler effects impossible.
@@
It is important to note that the use of near immunity to spoilers in our main result (Theorem 3.8) could be replaced by Condorcet consistency (Proposition 3.10). We use the framing in terms of spoilers for three reasons. First, near immunity to spoilers is strictly weaker than Condorcet consistency for three alternatives (while Condorcet consistency implies near immunity to spoilers on this domain, the converse implication fails, as shown by, e.g., the trivial voting method that selects all alternatives in every profile or by the Weighted Covering method mentioned above). Second, the framing in terms of immunity to spoilers highlights the desirability of using Minimax for elections with three alternatives, such as the top-three elections advocated in Foley and Maskin 2024, since mitigating spoiler effects is a significant benefit. Third, looking ahead to profiles with more than three alternatives (Section 5), Condorcet consistency does not imply near immunity to spoilers for more than three alternatives, and Minimax is one of the rare voting methods (along with Split Cycle) satisfying the axiom for any number of alternatives (see Table 1).
@@
    $F$ satisfies block preservation if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$, if $X(\mathbf{P})=X(\mathbf{P}^{\prime})$, $V(\mathbf{P})\cap V(\mathbf{P}^{\prime})=\varnothing$, and $\mathbf{P}^{\prime}$ contains for each linear order on $X(\mathbf{P})$ exactly one voter submitting that linear order and no other voters, then $\mathbf{P}+\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ and $F(\mathbf{P})\subseteq F(\mathbf{P}+\mathbf{P}^{\prime})$.